# Non-Resolution Reasoning (NRR): A Computational Framework for Contextual Identity and Ambiguity Preservation

### v5

@@ Abstract
Current artificial intelligence systems, despite remarkable capabilities in text generation and pattern recognition, exhibit a fundamental architectural limitation: they resolve ambiguity prematurely. This premature semantic collapse—the tendency to collapse multiple valid interpretations into a single output—stems from classical identity assumptions embedded in standard neural architectures. We propose Non-Resolution Reasoning (NRR), a computational framework that treats ambiguity retention as a valid reasoning mode rather than a defect to be eliminated. NRR introduces three core principles: (1) Non-Identity ($A\neq A$)—the same symbol refers to different entities across contexts; (2) Approximate Identity ($A\approx A$)—entities share partial structural overlap without being identical; and (3) Non-Resolution—conflicting interpretations can coexist without forced convergence. We formalize these principles through three architectural components: Multi-Vector Embeddings for context-dependent representation, Non-Collapsing Attention for parallel interpretation retention, and Contextual Identity Tracking (CIT) for maintaining $A\neq A$ across inference. We demonstrate NRR’s advantages through case studies in paradox handling, creative generation, and context-dependent reasoning. Crucially, we provide functional verification through a Turn 1 Entropy measurement demonstrating ambiguity preservation directly: NRR-lite maintains high entropy ($H=0.63$, near-maximum $0.69$) at the ambiguous turn, while standard architectures collapse early ($H=0.10$), demonstrating that NRR preserves interpretive flexibility until context arrives. NRR challenges the assumption that meaning must collapse to be useful, offering a foundation for AI systems capable of sophisticated ambiguity handling and creative reasoning. The question is not whether AI should resolve ambiguity, but when, how, and under whose control.
@@ 1.2 Contributions of This Work
    Functional Verification: We validate ambiguity preservation directly via Turn 1 Entropy measurement, showing NRR-lite maintains $H=0.63$ (near-maximum) while baselines collapse to $H=0.10$, robust across multiple random seeds.
@@
    Theoretical Formalization: We provide mathematical foundations, including operators for non-identity ($\neq$) and non-resolution ($\aleph$), with explicit formalization of the $A\neq A\approx A$ principle.
@@
    Implementation Roadmap: We provide concrete benchmarking strategies and computational cost analysis to enable validation and extension.
@@ 4 NRR in Practice: Case Studies
We illustrate how NRR represents internal states in scenarios challenging for conventional architectures.
@@ 4.1 Paradox Handling: Internal Representation Matters
Problem: “This sentence is false.” Standard systems attempt binary truth evaluation, leading to contradiction or infinite loops. NRR maintains multiple contextual interpretations (truth_context, falsity_context, paradox_context) in parallel as a stable internal structure, allowing reasoning to continue without forced resolution. The paradox becomes data rather than error.
@@ 4.2 Creative Generation: Polysemy in Writing
Task: “Write a story about ‘light’.” Conventional AI disambiguates early (e.g., to “illumination”), producing single-layer output. NRR retains multiple meanings (illumination, weight, mood) simultaneously during generation, enabling multi-layered, poetic output where meanings are interwoven.
@@ 4.3 Context-Dependent Identity
Problem: Context shift in dialogue (e.g., “The bank is solid” $\to$ “ducks are swimming”). Conventional AI must backtrack and reprocess when the initial assumption (financial) proves wrong. NRR continuously reweights coexisting identities (financial vs. river) without backtracking, as both interpretations were maintained from the start.
@@ 5 Implementation Framework
This section presents architectural components for implementing NRR.
@@ 5.1 Multi-Vector Embeddings
Unlike standard embeddings that collapse to single vectors, NRR embeddings retain multiple contextual variants simultaneously.
@@
These can be trained via contrastive initialization or adversarial diversity loss to ensure variants capture distinct semantic dimensions.
@@ 5.2 Non-Collapsing Attention
Standard attention uses softmax normalization, forcing probabilistic collapse. NRR uses non-collapsing functions (e.g., sigmoid) to allow multiple positions to receive strong attention simultaneously without competition.
@@ 4. Aggregate:
Stability is maintained via gradient clipping and attention strength regularization.
@@ 5.3.1 Context Detection
a new context ID is assigned. $\tau_{\text{context}}$ may be learned or fixed; in our minimal implementation, it is defined as the cosine-distance threshold separating intra-turn variability from inter-turn shifts.
@@ 5.4 Strategic Resolution: When and How to Collapse
While NRR treats non-resolution as computationally valid, practical systems must eventually act. We propose a resolution operator $\rho$ controlled by task affordances (classification vs. generation), contextual stability (context drift $\Delta c_{t}$), and interpretation dominance. Crucially, these policies are external to the NRR core, allowing the same representation to serve diverse tasks.
@@ 6 Functional Verification
Scope of this section: This section does not aim to establish empirical superiority of NRR over existing large language models. Instead, we provide a minimal functional verification (proof-of-concept) demonstrating that the proposed architecture behaves as intended: it preserves ambiguity before contextual cues arrive and resolves it afterward. Comparative evaluation against production LLMs is reserved for future work.
@@ 6.2 Model Comparison
  • NRR-lite: Maintains $k=2$ separate embeddings for the ambiguous token (“bank”). A context-dependent gate selects the appropriate embedding based on Turn 2 context. When Turn 2 is neutral (no context), the gate remains balanced.
Both models were trained on 1,000 samples with balanced labels (50% FINANCIAL, 50% RIVER) for 100 epochs.
@@ 6.3 Results
    Robustness: Results replicate across 5 random seeds ($t=13.3$, $p<0.001$).
This verification confirms that NRR-lite implements the $A\neq A$ principle: the same symbol (“bank”) maintains distinct potential identities until context determines which applies.
@@ 8 Conclusion
Theoretical foundations and philosophical framework will be addressed in a forthcoming companion paper. This paper introduced Non-Resolution Reasoning (NRR), a framework challenging classical identity assumptions in AI. We demonstrated through theory and functional verification that preserving ambiguity is not a cost but a capability. Our Turn 1 Entropy verification showed that NRR-lite maintains high interpretive entropy ($H=0.63$) when context is absent, while standard architectures collapse prematurely (H=0.10H=0.10)—yet both achieve perfect accuracy when context arrives. NRR challenges the assumption that meaning must collapse to be useful. The question is not whether AI should resolve ambiguity, but when, how, and under whose control.

### v6

@@ Abstract
Current artificial intelligence systems, despite remarkable capabilities in text generation and pattern recognition, exhibit a fundamental architectural limitation: they resolve ambiguity prematurely. This premature semantic collapse—the tendency to collapse multiple valid interpretations into a single output—stems from classical identity assumptions embedded in standard neural architectures. We propose Non-Resolution Reasoning (NRR), a computational framework that treats ambiguity retention as a valid reasoning mode rather than a defect to be eliminated. NRR introduces three core principles: (1) Non-Identity ($A\neq A$)—the same symbol refers to different entities across contexts; (2) Approximate Identity ($A\approx A$)—entities share partial structural overlap without being identical; and (3) Non-Resolution—conflicting interpretations can coexist without forced convergence. We formalize these principles through three architectural components: Multi-Vector Embeddings for context-dependent representation, Non-Collapsing Attention for parallel interpretation retention, and Contextual Identity Tracking (CIT) for maintaining $A\neq A$ across inference. We illustrate NRR’s potential through case studies in paradox handling, creative generation, and context-dependent reasoning. Crucially, we provide functional verification in a synthetic two-turn disambiguation task through a Turn 1 Entropy measurement demonstrating ambiguity preservation directly: NRR-lite maintains high entropy ($H=0.63$, near-maximum $0.69$) at the ambiguous turn, while standard architectures collapse early ($H=0.10$), demonstrating that NRR preserves interpretive flexibility until context arrives. NRR challenges the assumption that meaning must collapse to be useful, offering a foundation for AI systems capable of sophisticated ambiguity handling and creative reasoning. The question is not whether AI should resolve ambiguity, but when, how, and under whose control.
@@ 1.2 Contributions of This Work
    Functional Verification: We validate ambiguity preservation directly via Turn 1 Entropy measurement, showing NRR-lite maintains $H=0.63$ (near-maximum) while the baseline collapses to $H=0.10$, consistently across five random seeds.
@@
    Theoretical Formalization: We provide a mathematical outline, including a formal state space and eight operators for non-collapsing computation, with explicit formalization of the $A\neq A\approx A$ principle.
@@
    Implementation Roadmap: We provide concrete benchmarking strategies and computational cost considerations to enable validation and extension.
@@ 4 NRR in Practice: Case Studies
We illustrate how NRR would represent internal states in scenarios challenging for conventional architectures. These are conceptual illustrations; empirical validation is provided in Section 6.
@@ 4.1 Paradox Handling: Internal Representation Matters
Problem: “This sentence is false.” Standard systems attempt binary truth evaluation, leading to contradiction or infinite loops. NRR would maintain multiple contextual interpretations (truth_context, falsity_context, paradox_context) in parallel as a stable internal structure, allowing reasoning to continue without forced resolution. The paradox becomes data rather than error.
@@ 4.2 Creative Generation: Polysemy in Writing
Task: “Write a story about ‘light’.” Conventional AI disambiguates early (e.g., to “illumination”), producing single-layer output. NRR would retain multiple meanings (illumination, weight, mood) simultaneously during generation, enabling multi-layered, poetic output where meanings are interwoven.
@@ 4.3 Context-Dependent Identity
Problem: Context shift in dialogue (e.g., “The bank is solid” $\to$ “ducks are swimming”). Conventional AI must backtrack and reprocess when the initial assumption (financial) proves wrong. NRR would continuously reweight coexisting identities (financial vs. river) without backtracking, as both interpretations were maintained from the start.
@@ 5 Implementation Framework
This section presents architectural components for implementing NRR. While related to prior work on multi-prototype embeddings and non-softmax attention, the mechanisms in NRR are unified toward explicit ambiguity preservation rather than sense selection or sparsity.
@@ 5.1 Multi-Vector Embeddings
Unlike standard embeddings that collapse to single vectors, NRR embeddings are designed to retain multiple contextual variants simultaneously.
@@
These could be trained via contrastive initialization or adversarial diversity loss to ensure variants capture distinct semantic dimensions; we leave concrete training schemes to future empirical work.
@@ 5.2 Non-Collapsing Attention
Standard attention uses softmax normalization, forcing probabilistic collapse. NRR proposes non-collapsing functions (e.g., sigmoid) to allow multiple positions to receive strong attention simultaneously without competition.
@@ 4. Aggregate:
Stability could be maintained via gradient clipping and attention strength regularization.
@@ 5.3.1 Context Detection
a new context ID is assigned. $\tau_{\text{context}}$ may be learned or fixed; in a minimal implementation, it could be defined as the cosine-distance threshold separating intra-turn variability from inter-turn shifts.
@@ 5.4 Strategic Resolution: When and How to Collapse
While NRR treats non-resolution as computationally valid, practical systems must eventually act. We propose a resolution operator controlled by task affordances (classification vs. generation), contextual stability (context drift $\Delta c_{t}$), and interpretation dominance. Crucially, these policies are external to the NRR core, allowing the same representation to serve diverse tasks.
5.5 Protocol Extensions: State Space and Operators
The architectural components above (MVE, NCA, CIT) can be unified under a formal state space that enables systematic extension. This section provides the minimal specification for researchers implementing NRR-based systems.
5.5.1 State Space Formalization
The non-collapsing state is represented as:
$\mathcal{S}=\bigl((v_{i},c_{i},w_{i},m_{i})\bigr)_{i=1}^{n}$ (4)
where $v_{i}\in\mathbb{R}^{d}$ is a semantic vector, $c_{i}$ is a context identifier, $w_{i}\in[0,1]$ is an activation weight, and $m_{i}$ is optional metadata. Crucially, the weights $w_{i}$ are not normalized to sum to unity—multiple interpretations can maintain high activation simultaneously without competing for probability mass. This is the formal expression of non-collapse.
For measurement purposes, we associate a discrete distribution over interpretations:
$p(i\mid\mathcal{S})=\frac{w_{i}}{\sum_{j}w_{j}}$ (5)
and define $H(\mathcal{S})$ as the Shannon entropy of this distribution:
$H(\mathcal{S})=-\sum_{i=1}^{n}p(i\mid\mathcal{S})\log p(i\mid\mathcal{S})$ (6)
This normalization is used only for evaluation; the internal NRR dynamics do not require $w_{i}$ to sum to unity.
5.5.2 Interference Operators
We introduce eight candidate interference operators that act on $\mathcal{S}$. We define an NRR interference operator as a transformation $T$ on $\mathcal{S}$ whose intended property is approximate information preservation, i.e., our design goal is that
$H(T(\mathcal{S}))\geq H(\mathcal{S})-\epsilon$ (7)
for small $\epsilon$. Table 1 specifies the operators and their roles; verifying and refining their information-preservation guarantees is left to future empirical and theoretical work. One operator, CPP Integration ($\kappa$), implements the Contradiction-Preservation Principle: contradictory hypotheses are retained rather than eliminated.
Operator Symbol Function
Stripping $\sigma$ Removes learned biases inducing collapse
Label-Free Abstraction $\alpha$ Projects to geometric relations without labels
Positioning $\rho$ Assigns contextual coordinates (implements CIT)
Invariance $\iota$ Identifies structure stable across contexts
Dampening $\delta$ Reduces activation differentials
Deferred Resolution $\tau$ Postpones projection to output boundary
CPP Integration $\kappa$ Preserves contradictory states with conflict tags
Persistence $\pi$ Maintains state across turns with decay
Table 1: NRR interference operators. Each is intended to transform $\mathcal{S}$ in ways that avoid collapse.
A typical pipeline: $\mathcal{S}^{\prime}=\pi(\kappa(\delta(\rho(\sigma(\mathcal{S}))),\mathcal{S}_{\text{new}}))$—stripping biases, positioning, dampening, integrating new information, and persisting.
5.5.3 Non-Destructive Projection
To produce singular output while preserving internal state, NRR uses non-destructive projection:
$\Pi(\mathcal{S})=g(v_{i^{*}}),\quad i^{*}=\arg\max_{i}w_{i},\quad\mathcal{S}^{\prime}=\mathcal{S}$ (8)
where $g:\mathbb{R}^{d}\to\mathcal{Y}$ maps the semantic vector to the output space (e.g., token vocabulary). The system outputs $g(v_{i^{*}})$ but retains the full state $\mathcal{S}$. This contrasts with collapse-based systems where output destroys alternatives. The user receives a singular answer; the system preserves interpretive depth for subsequent reasoning.
@@ 6 Functional Verification
Scope of this section: This section provides a controlled synthetic evaluation demonstrating that the proposed architecture behaves as intended: it preserves ambiguity before contextual cues arrive and resolves it afterward. This is not a comparison against modern LLMs but a minimal functional verification (proof-of-concept). Comparative evaluation against production systems is reserved for future work.
@@ 6.2 Model Comparison
  • NRR-lite: Implements the MVE principle in minimal form, maintaining $k=2$ separate embeddings for the ambiguous token (“bank”). A context-dependent gate selects the appropriate embedding based on Turn 2 context. When Turn 2 is neutral (no context), the gate remains balanced.
Both models were trained on 1,000 samples with balanced labels (50% FINANCIAL, 50% RIVER) for 100 epochs. All reported entropy values and statistics are computed from the trained models over 5 independent random seeds.
@@ 6.3 Results
    Robustness: Results replicate across 5 random seeds (paired $t$-test: $t(4)=13.3$, $p<0.001$).
This verification provides evidence that NRR-lite implements the $A\neq A$ principle: the same symbol (“bank”) maintains distinct potential identities until context determines which applies.
@@ 8 Conclusion
This paper introduced Non-Resolution Reasoning (NRR), a framework challenging classical identity assumptions in AI. In our controlled setting, we showed that preserving ambiguity can function as a capability rather than a cost. Our Turn 1 Entropy verification showed that NRR-lite maintains high interpretive entropy ($H=0.63$) when context is absent, while standard architectures collapse prematurely (H=0.10H=0.10)—yet both achieve perfect accuracy when context arrives. NRR challenges the assumption that meaning must collapse to be useful. The question is not whether AI should resolve ambiguity, but when, how, and under whose control.